OpenStudy (rootbeer003):

What is the formula for calculating the length of the space diagonal of a right rectangular prism on the basis of the prism's length l, width w, and height h?

OpenStudy (kg1975):

This is Pythagorean theorem for a 3D Triangle basically... A^2+B^2+C^2 = D^2 edit: all rectangular prisms are "right angled" ;D

OpenStudy (mathmale):

rootie: Haven't we discussed a similar question before? If I recall correctly, the previous question has actual, numerical measurements shown, whereas now you have only l, w and h. Could you not apply what you learned in the previous problem to develop the desired formula here?

OpenStudy (mathmale):

Check out the formula given you by kg1975. Can you derive his / her formula?

OpenStudy (kg1975):

her xD

OpenStudy (rootbeer003):

Sorry but those arnt one of the choices..

OpenStudy (kg1975):

Option A: d = l2 + w2 + h2 Option B: d = sqrt l2 + w2 + h2 Option C: d= h2 = l2 + w2 Option D: d =square rt l + w +h are these your choices?

OpenStudy (rootbeer003):

sort of

OpenStudy (kg1975):

u can use the formula i gave u to make another formula thats the answer. ive had this questuon before... the answer is B

OpenStudy (kg1975):

ok?

OpenStudy (rootbeer003):

O k but im going to look into it myself

OpenStudy (kg1975):

mmkay